# Fairshare Data Pricing via Data Valuation for Large Language Models

### v2

@@ 4.3 Ablation Study: Effect of Different Data Valuation Methods
Setup. Using the models and datasets introduced earlier in Section 4.1, we run separate simulations for each market (math, medical, physical reasoning), testing different data valuation methods for the buyer. Each buyer is randomly assigned to use one of four valuation methods to select training data and fine-tune their model. The seller receives payments according to the data’s assigned value. Valuation scores are normalized to $[0,1]$ for pricing compatibility.
@@
Data Valuation Methods. We consider the following methods, where each assigns a value to every data sample: (1) Constant baseline – assigns the same value, mimicking flat-rate pricing on platforms; (2) Random baseline – values drawn uniformly from $[0,1]$; (3) Semantic – uses BM25 [30] to compute average similarity to the representative set; (4) Influence-based: returns a score which leverages learning gradients to estimate a data sample’s avg. contribution to model learning of a representative dataset. Specifically, we use $\text{Infl}_{\text{IP}}$ [56, 22] and DataInf  [21] (see Appendix A). In addition, we provide experiments using data valuation through model re-training, which is considered an “oracle" for influence-based methods [23, 56] in Section F.2. Further details on the experiment (e.g., model training) are described in Section E.1.
@@
Results. Figure 6 presents results across data valuation methods. Buyers using a valuation method (ie..g, BM25, $\text{Infl}_{\text{IP}}$, DataInf) in general achieved higher model performance across tasks. When considering the trade-off between cost and performance, $\text{Infl}_{\text{IP}}$ offered the best balance – delivering strong model improvements at a lower cost than constant, random, and BM25 (see Figures 6(a) and 9(a)).
Our results highlight the benefits of learning-aware data valuation methods. By prioritizing high-impact data, they offer a better alternative for buyers, particularly those with limited budgets.
@@
We also run the error analysis of data valuation methods. Following previous studies in approximating true data value [62, 63], we compute the correlation between oracle and $\text{Infl}_{\text{IP}}$, reporting spearman correlation of $0.54,0.42$ for MathQA and PIQA (see Figure 8 in Appendix F). As previously noted, our framework is method-agnostic and can directly benefit from future advances in data valuation accuracy and scalability.
@@
Figure 6: Left column: Average price-per-sample cost of purchased data across math, medical, and physical reasoning data markets using different data valuation methods. Right, middle-columns: Buyers’ model performance versus cost, before purchasing (BP) data, and after purchasing 2K and 4K data samples. Additional analysis on the Pythia-1b/410m models are in Appendix F.
@@ 5 Conclusion
We introduced a fairshare pricing framework that leverages data valuation methods to enable transparent, equitable pricing for LLM training data. Our results show that buyers achieve greater performance at lower cost, while sellers receive optimal compensation – yielding a win-win outcome that promotes long-term market sustainability. To our knowledge, this is the first work integrating game-theoretic pricing with LLM-specific data valuation, capturing real-world dynamics of the emerging data market. The framework offers actionable guidance for policymakers aiming to ensure fairness and transparency, while empowering small firms and startups through fair market access. Future work could extend this framework to support additional data valuation methods, incomplete information settings (e.g., Bayesian games), and broader data domains (e.g., pretraining vs. fine-tuning). We hope this work advances equitable market design in AI and emerging technologies.
@@ Appendix A Influence-based Data Valuation
$\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})$ (12)
Instead of taking a single training step to measure the influence of $d\in D$ on $d^{\prime}$, we can instead approximate Equation 12 with using the following:
@@ Appendix B Royalty model
So far, we have shown the case of flat rate (see Section 3.1), which is well-suited resource-rich buyers, such as leading tech companies whose LLMs generate significant economic value due to their wide-ranging impact and scalability. In this section, we introduce the royalty model, a contract framework that differs from the flat rate by offering a subscription-like structure. Under the royalty model, the price paid for training data is proportional to the future economic value generated by the LLM, providing a flexible and performance-based approach to data valuation. This scenario incorporates buyers in a less dominant position – those who are (1) uncertain about the prospective model outcome or (2) do not own a sufficient cash flow for purchasing data with full prices. We present updated decision-making models for buyers and sellers as follows.
@@
On the other hand, similar to the budget constraint (see Equation 2), here each buyer $B_{k}$ has a maximum rate $\overline{\alpha}_{k}$ that it is willing to pay. Then the buyer’s purchasing problem is given as
[…]
$\mathcal{X}_{k,N,\text{frac}}$ :=$:=\{\boldsymbol{x}\mid g_{k,N,\text{frac}}(\boldsymbol{x})\geq 0,f(\boldsymbol{\alpha},\boldsymbol{x})\leq\overline{\alpha}_{k}\},$ (16)
@@
[…]
$\mathcal{A}_{j,M}$ :=$:=\{{\alpha}_{j}\in[0,1)\mid r_{\text{frac}}(\alpha_{j})\geq 0\},$ (19)
@@ B.1 Solving for optimal price of the royalty model
Solve seller’s problem. First, we consider when $\alpha_{j}=0$. We could first find all $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}$ such that $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$. And we denote the set that contains such $\boldsymbol{x}$ as $\mathcal{X}_{k,N-1,\text{frac}}^{1}$. If $\mathcal{X}_{k,N-1,\text{frac}}^{1}$ is empty, then $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}=0$ (indicating whether $B_{k}$ will purchase $D_{j}$ at price $p_{j}$ or not), as $S_{j}$ cannot bring positive value to $B_{k}$; else, then for all $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, thanks to the monotonicity of $f_{j}$ over $\alpha_{j}$, we could gradually increase $\alpha_{j}$ until the either of the two criterion are met first: (1) we find the largest $\alpha_{j}$ such that $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$, and (2) $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})\leq\overline{\alpha}_{k}$. Then we have the following property about the optimal rate $\alpha_{j}^{*}$ for Equation 18:
@@
Observing from Lemmas 5 and 2, we see that the both the optimal price $p^{*}_{j}$ and the optimal rate $\alpha_{j}^{*}$ are closely tied to $B_{k}$’s maximum willingness to pay. That is, compared to the market prior to the arrival of $S_{j}$, the optimal values are characterized by the minimum of two factors: (1) marginal utility that $S_{j}$ provides to $B_{k}$ and (2) $B_{k}$’s budget surplus. It is also noted that, under royalty model, the rate function $f$ also plays an important role as it determines the how the single rate $\alpha_{j}$ affects the total rate that $B_{k}$ pays.
@@ Appendix C Applications for Real-Life Scenarios
In real-life settings, the relationship between the data valuation of a training sample and the buyer’s utility $u_{k}$ (i.e., the economical value, which may be expressed in dollar amounts) can have different mappings, as mentioned in Section 3.1. Suppose the data valuation function is denoted as $v_{k}:D\rightarrow\mathbb{R}$ for a dataset $D$. Then, a buyer may expect a linear relationship between $v_{k}$ and $u_{k}$, where the utility increases as the data valuation score increases. Alternatively, a buyer may prefer to only purchase data beyond a certain threshold for $v_{k}$. In this section, we present three types mappings between $v_{k}$ and $u_{k}$ to reflect these scenarios: linear, discrete, and zero-one mappings. We show that these mappings can be easily adapted to our proposed framework in Section 3. We only present the updated buyer’s purchasing problem (Equation 2) since the seller’s pricing problem (Equation 4) stays the same.
@@ C.1 Linear Outcome
In practice, there are many applications where $u_{k}$ is an affine function of $v_{k}$. As previously mentioned, training LLMs on data with higher valuation scores $v_{k}$ can result in better economic value towards downstream model performance, as shown in previous works [22, 23]. In this outcome setting, in addition to considering $u_{k}$ to be a affine function of $v_{k}$, we also include a bias variable $\beta$ to account for other potential other factors that are independent of $v_{k}$. Therefore, we can set $u_{k}=\gamma v_{k}(\boldsymbol{x})+\beta$ into Equation 1, where $\gamma\in\mathbb{R}_{+}$ is a known coefficient, and obtain buyer $B_{k}$’s net utility function for the linear outcome:
@@
To obtain optimal price $p^{*}$, we can directly refer to same procedure described in Section 3 using set values for $\gamma$ and $\beta$.
@@ C.2 Discrete Outcome
We again apply the same procedure in Section 3 to solve for the optimal pricing.
@@ C.3 Zero-One Outcome
which is an affine function of $v_{k}$. Therefore, we again apply same procedure in Section 3 to solve for the optimal pricing.
@@ C.4 Multiple tasks
whose solution could adopt the same procedure as described in Section 3 to solve for the optimal pricing.
@@ Appendix D Lemmas and Proofs
Lemma 1. With Assumptions 1, 1.1 and 2, any exploitative pricing (i.e., $p_{t}<p_{t}^{*},\forall t$) will only maximize cumulative utility within a finite horizon – after which it is strictly suboptimal.
@@
Lemma 1 equivalently states that with Assumptions 1, 1.1 and 2, the optimal pricing strategy for the buyer is also the ideal price $p^{*}_{t}$. Thus, we show that when the buyer pays the ideal price $p^{*}_{t}$, its total value is the largest, i.e.,
@@
Therefore, we must have $G\geq\frac{\min_{t\in[0,\infty)}\mathbb{E}[u_{t}-p^{*}_{t}]}{1-\delta}$. Along with Assumptions 1, 1.1 and 2, this gives us, for a given $u_{t}$ and $b_{t}$,
@@
With Assumption 1, let $P:=\lim_{T\rightarrow\infty}P_{T}$ and $S:=\sum_{i=0}^{\infty}(1-\pi(p_{i},p^{*}_{i}))$. Then for the class of all exploitative pricing strategies where $p_{t}\leq p_{t}^{*},\forall t$, we have (1) $P_{T}$ is strictly decreasing and (2) The limit of reduced participation is lower bounded, i.e., $01-P\geq 1-e^{-S}>$ 0.
@@
The first part is trivial to see as we assume that $\pi(p_{i},p^{*}_{i})<1$ for all $p_{i}<p^{*}_{i}$ from Assumption 1.
@@
Since we have $10<\pi(p_{i},p^{*}_{i})<$ 1, then $\log\left(\pi(p_{i},p^{*}_{i})\right)\leq\pi(p_{i},p^{*}_{i})-1$, indicating that
@@
Lemma 2. Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP:
@@
Recall that without dataset $D_{j}$, each buyer $B_{k}$ has already solved $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}g_{k,N-1}(\boldsymbol{x})$ according to our market definition in Section 3, where $\mathcal{X}_{k,N-1}$ is the set of all feasible purchase decisions. Next, after seller $S_{j}$ (with $D_{j}$) has arrived on the market, we analyze the conditions in which $B_{k}$ will purchase $D_{j}$ at a potential price $p_{j}$. For each feasible purchase decision (i.e., a collection of datasets), represented by $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$, let $\boldsymbol{x}+e_{j}$ denote its union with $D_{j}$, where $e_{j}$ is the unit vector indicating $D_{j}$ is selected. For buyer $B_{k}$ to change their previous decision to purchase $D_{j}$, there are two requirements that need to be satisfied. First, we must have:
@@
which ensures that purchasing $D_{j}$ does not exceed the buyer’s budget $b_{k}$. If both requirements are satisfied, then the buyer $B_{k}$ will change their previous purchasing decision in order to purchase $D_{j}$ under the price $p_{j}$. This procedure is presented in detail in Algorithm 1 in Section F.1.
@@
Lemma 3 The optimal price for the seller $S$ under our framework is
@@
With assumptions 1, 1.1 and 2, $p^{*}_{t}$ gives the buyer the maximum cumulative net utility over infinite horizon.
@@
Further, we refer to the proof of Lemma 1 for the second part. ∎
Lemma 4. Suppose we have a LLM with parameters $\theta$. We perform a gradient descent step with training sample $d$ with learning rate $\eta$ such that $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Then,
@@
$\mathcal{L}(d^{\prime};\hat{\theta})$ $=\mathcal{L}(d^{\prime};\theta)+\nabla\mathcal{L}(d^{\prime};\theta)\dot{(}\hat{\theta}-\theta)+\mathcal{O}(||\hat{\theta}-\theta||^{2})$ (37)
$\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})$ $=-\nabla\mathcal{L}(d^{\prime};\theta)\dot{(}\hat{\theta}-\theta)+\mathcal{O}(||\hat{\theta}-\theta||^{2})$ (38)
@@
$\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})$ $\approx\eta\nabla\mathcal{L}(d^{\prime};\theta)\cdot\nabla\mathcal{L}(z;\theta)\propto\nabla\mathcal{L}(d^{\prime};\theta)\cdot\nabla\mathcal{L}(d;\theta)$ (39)
@@
Lemma 5 Define $\alpha_{j}^{\boldsymbol{x}}$ as
@@
Lemma 7 (The Trade-Off Threshold Is Increasing as $\delta$ Decreases).
And $t^{*}$ is increasing as $\delta$ increases.
Proof.

### v3

@@ 4.3 Ablation Study: Effect of Different Data Valuation Methods
Setup. Using the models and datasets introduced earlier in Section˜4.1, we run separate simulations for each market (math, medical, physical reasoning), testing different data valuation methods for the buyer. Each buyer is randomly assigned to use one of four valuation methods to select training data and fine-tune their model. The seller receives payments according to the data’s assigned value. Valuation scores are normalized to $[0,1]$ for pricing compatibility.
@@
Data Valuation Methods. We consider the following methods, where each assigns a value to every data sample: (1) Constant baseline – assigns the same value, mimicking flat-rate pricing on platforms; (2) Random baseline – values drawn uniformly from $[0,1]$; (3) Semantic – uses BM25 [30] to compute average similarity to the representative set; (4) Influence-based: returns a score which leverages learning gradients to estimate a data sample’s avg. contribution to model learning of a representative dataset. Specifically, we use $\text{Infl}_{\text{IP}}$ [56, 22] and DataInf  [21], which are influence-based methods adapted for the LLM realm. See Appendix˜A for additional details.
Furthermore, we adopt a one-step training approach for data valuation, where the value of each data sample is estimated by performing a single training step on it and measuring the resulting change in model performance relative to the original model. Results are presented in in Section˜F.2. This approach serves as an “oracle” for influence-based methods [23, 56], as it directly quantifies data value through model training, although it is more computationally expensive than the previous listed methods.
Market/Pricing Setup: We reserve $1\%$ of the samples from each dataset’s training split to represent the existing data in their respective markets. Each data sample was randomly priced between $(0,1]$. Next, for each remaining data sample in the training set, we determine whether each buyer will purchase the data sample at potential price points $[0.5,0.625,0.75,0.875,1.0]$ by solving Equation˜2. The seller then sets their prices according to Equation˜4. We price data separately for each data valuation method. This assesses the method’s ability to discern whether a new data sample is worth purchasing for each buyer given the existing market data, as noted in our analysis in Section˜3. Further details on all experiments are described in Section˜E.1.
@@
Results. Figure˜6 presents results across data valuation methods. Buyers using a valuation method (ie., BM25, $\text{Infl}_{\text{IP}}$, DataInf) in general achieved higher model performance across tasks. When considering the trade-off between cost and performance, $\text{Infl}_{\text{IP}}$ offered the best balance, delivering strong model improvements at a lower cost than constant, random, and BM25 (Figures 6(a) and 9(a)). Our results highlight the benefits of learning-aware data valuation methods. By prioritizing high-impact data, they offer a better alternative for buyers, particularly those with limited budgets.
@@
We also run the error analysis of data valuation methods. Following previous studies [71, 72], we compute the correlation between Oracle and $\text{Infl}_{\text{IP}}$, reporting spearman correlation of $0.54,0.42$ for MathQA and PIQA (see Figure˜8 in Appendix˜F). This underscores the opportunity for future advances in data valuation accuracy and scalability, which can be seamlessly integrated into our flexible fairshare framework.
@@
Figure 6: Left column: Average price-per-sample cost of purchased data across math, medical, and physical reasoning data markets using different data valuation methods. Right, middle-columns: Buyers’ model performance versus cost, before purchasing (BP) data, and after purchasing 2K and 4K data samples. Additional analysis on the Pythia-1b/410m models are in Appendix˜F.
@@ 5 Conclusion
We introduced a fairshare pricing framework that leverages data valuation methods to enable transparent and equitable training data pricing for LLMs. Our results show that buyers achieve higher gains at lower cost. At the same time, sellers earn optimal prices for their data, fostering a win-win outcome that enhances long-term market sustainability and social welfare. To our knowledge, this is the first work to integrate game-theoretic pricing models with LLM-specific data valuation techniques, addressing the real-world dynamics of the emerging data market. Our framework offers actionable insights for policymakers and regulators aiming to ensure fairness and transparency in LLM training data markets. By fostering fair market access, our framework also empowers small businesses and startups, leading to more equitable technological advancements. Future work could extend our framework to incorporate additional data valuation methods, incomplete information settings (e.g., Bayesian games), and diverse data domains (e.g., pretraining vs. fine-tuning). We hope this work paves a fruitful way for future research in equitable markets for AI and emerging technologies.
@@ Appendix A Influence-based Data Valuation
[…]
Instead of taking a single training step to measure the influence of $d\in D$ on $d^{\prime}$, we can instead approximate Equation˜13 with using the following:
@@ Appendix B Royalty model
So far, we have shown the case of flat rate (see Section˜3.1), which is well-suited resource-rich buyers, such as leading tech companies whose LLMs generate significant economic value due to their wide-ranging impact and scalability. In this section, we introduce the royalty model, a contract framework that differs from the flat rate by offering a subscription-like structure. Under the royalty model, the price paid for training data is proportional to the future economic value generated by the LLM, providing a flexible and performance-based approach to data valuation. This scenario incorporates buyers in a less dominant position – those who are (1) uncertain about the prospective model outcome or (2) do not own a sufficient cash flow for purchasing data with full prices. We present updated decision-making models for buyers and sellers as follows.
@@
On the other hand, similar to the budget constraint (see Equation˜2), here each buyer $B_{k}$ has a maximum rate $\overline{\alpha}_{k}$ that it is willing to pay. Then the buyer’s purchasing problem is given as
$\tilde{\boldsymbol{x}}^{k,N,\text{frac}}$ $:=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}_{k,N,\text{frac}}}\;g_{k,N,\text{frac}}(\boldsymbol{x}),\quad\text{s.t.}$ (16)
$\mathcal{X}_{k,N,\text{frac}}$ $:=\{\boldsymbol{x}\mid g_{k,N,\text{frac}}(\boldsymbol{x})\geq 0,f(\boldsymbol{\alpha},\boldsymbol{x})\leq\overline{\alpha}_{k}\},$ (17)
@@
$\alpha^{*}_{j}$ $:=\operatorname*{arg\,max}_{{\alpha}_{j}\in\mathcal{A}_{j,M,\text{frac}}}\;r_{\text{frac}}(\alpha_{j}),\quad\text{s.t.}$ (19)
$\mathcal{A}_{j,M}$ $:=\{{\alpha}_{j}\in[0,1)\mid r_{\text{frac}}(\alpha_{j})\geq 0\},$ (20)
@@ B.1 Solving for optimal price of the royalty model
Solve seller’s problem. First, we consider when $\alpha_{j}=0$. We could first find all $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}$ such that $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$. And we denote the set that contains such $\boldsymbol{x}$ as $\mathcal{X}_{k,N-1,\text{frac}}^{1}$. If $\mathcal{X}_{k,N-1,\text{frac}}^{1}$ is empty, then $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}=0$ (indicating whether $B_{k}$ will purchase $D_{j}$ at price $p_{j}$ or not), as $S_{j}$ cannot bring positive value to $B_{k}$; else, then for all $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, thanks to the monotonicity of $f_{j}$ over $\alpha_{j}$, we could gradually increase $\alpha_{j}$ until the either of the two criterion are met first: (1) we find the largest $\alpha_{j}$ such that $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$, and (2) $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})\leq\overline{\alpha}_{k}$. Then we have the following property about the optimal rate $\alpha_{j}^{*}$ for Equation˜19:
@@
Observing from Lemmas˜6 and 2, we see that the both the optimal price $p^{*}_{j}$ and the optimal rate $\alpha_{j}^{*}$ are closely tied to $B_{k}$’s maximum willingness to pay. That is, compared to the market prior to the arrival of $S_{j}$, the optimal values are characterized by the minimum of two factors: (1) marginal utility that $S_{j}$ provides to $B_{k}$ and (2) $B_{k}$’s budget surplus. It is also noted that, under royalty model, the rate function $f$ also plays an important role as it determines the how the single rate $\alpha_{j}$ affects the total rate that $B_{k}$ pays.
@@ Appendix C Applications for Real-Life Scenarios
In real-life settings, the relationship between the data valuation of a training sample and the buyer’s utility $u_{k}$ (i.e., the economical value, which may be expressed in dollar amounts) can have different mappings, as mentioned in Section˜3.1. Suppose the data valuation function is denoted as $v_{k}:D\rightarrow\mathbb{R}$ for a dataset $D$. Then, a buyer may expect a linear relationship between $v_{k}$ and $u_{k}$, where the utility increases as the data valuation score increases. Alternatively, a buyer may prefer to only purchase data beyond a certain threshold for $v_{k}$. In this section, we present three types mappings between $v_{k}$ and $u_{k}$ to reflect these scenarios: linear, discrete, and zero-one mappings. We show that these mappings can be easily adapted to our proposed framework in Section˜3. We only present the updated buyer’s purchasing problem (Equation˜2) since the seller’s pricing problem (Equation˜4) stays the same.
@@ C.1 Linear Outcome
In practice, there are many applications where $u_{k}$ is an affine function of $v_{k}$. As previously mentioned, training LLMs on data with higher valuation scores $v_{k}$ can result in better economic value towards downstream model performance, as shown in previous works [22, 23]. In this outcome setting, in addition to considering $u_{k}$ to be a affine function of $v_{k}$, we also include a bias variable $\beta$ to account for other potential other factors that are independent of $v_{k}$. Therefore, we can set $u_{k}=\gamma v_{k}(\boldsymbol{x})+\beta$ into Equation˜1, where $\gamma\in\mathbb{R}_{+}$ is a known coefficient, and obtain buyer $B_{k}$’s net utility function for the linear outcome:
@@
To obtain optimal price $p^{*}$, we can directly refer to same procedure described in Section˜3 using set values for $\gamma$ and $\beta$.
@@ C.2 Discrete Outcome
We again apply the same procedure in Section˜3 to solve for the optimal pricing.
@@ C.3 Zero-One Outcome
which is an affine function of $v_{k}$. Therefore, we again apply same procedure in Section˜3 to solve for the optimal pricing.
@@ C.4 Multiple tasks
whose solution could adopt the same procedure as described in Section˜3 to solve for the optimal pricing.
@@ Appendix D Lemmas and Proofs
Lemma˜1. With Assumptions˜1, 1.1 and 2, any exploitative pricing (i.e., $p_{t}<p_{t}^{*},\forall t$) will only maximize cumulative utility within a finite horizon – after which it is strictly suboptimal.
@@
Lemma˜1 equivalently states that with Assumptions˜1, 1.1 and 2, the optimal pricing strategy for the buyer is also the ideal price $p^{*}_{t}$. Thus, we show that when the buyer pays the ideal price $p^{*}_{t}$, its total value is the largest, i.e.,
@@
Therefore, we must have $G\geq\frac{\min_{t\in[0,\infty)}\mathbb{E}[u_{t}-p^{*}_{t}]}{1-\delta}$. Along with Assumptions˜1, 1.1 and 2, this gives us, for a given $u_{t}$ and $b_{t}$,
@@
With Assumption˜1, let $P:=\lim_{T\rightarrow\infty}P_{T}$ and $S:=\sum_{i=0}^{\infty}(1-\pi(p_{i},p^{*}_{i}))$. Then for the class of all exploitative pricing strategies where $p_{t}\leq p_{t}^{*},\forall t$, we have (1) $P_{T}$ is strictly decreasing and (2) The limit of reduced participation is lower bounded, i.e., $1-P\geq 1-e^{-S}>0$.
@@
The first part is trivial to see as we assume that $\pi(p_{i},p^{*}_{i})<1$ for all $p_{i}<p^{*}_{i}$ from Assumption˜1.
@@
Since we have $0<\pi(p_{i},p^{*}_{i})<1$, then $\log\left(\pi(p_{i},p^{*}_{i})\right)\leq\pi(p_{i},p^{*}_{i})-1$, indicating that
@@
Lemma˜2. Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP:
@@
Recall that without dataset $D_{j}$, each buyer $B_{k}$ has already solved $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}g_{k,N-1}(\boldsymbol{x})$ according to our market definition in Section˜3, where $\mathcal{X}_{k,N-1}$ is the set of all feasible purchase decisions. Next, after seller $S_{j}$ (with $D_{j}$) has arrived on the market, we analyze the conditions in which $B_{k}$ will purchase $D_{j}$ at a potential price $p_{j}$. For each feasible purchase decision (i.e., a collection of datasets), represented by $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$, let $\boldsymbol{x}+e_{j}$ denote its union with $D_{j}$, where $e_{j}$ is the unit vector indicating $D_{j}$ is selected. For buyer $B_{k}$ to change their previous decision to purchase $D_{j}$, there are two requirements that need to be satisfied. First, we must have:
@@
which ensures that purchasing $D_{j}$ does not exceed the buyer’s budget $b_{k}$. If both requirements are satisfied, then the buyer $B_{k}$ will change their previous purchasing decision in order to purchase $D_{j}$ under the price $p_{j}$. This procedure is presented in detail in Algorithm˜1 in Section˜F.1.
@@
Lemma˜3 The optimal price for the seller $S$ under our framework is
@@
With assumptions˜1, 1.1 and 2, $p^{*}_{t}$ gives the buyer the maximum cumulative net utility over infinite horizon.
@@
Further, we refer to the proof of Lemma˜1 for the second part. ∎
Lemma˜4 The threshold time period where the fairshare pricing obtains higher cumulative utility than any class of exploitative pricing is:
And $t^{*}$ is increasing as $\delta$ increases.
Proof.
Lemma˜5. Suppose we have a LLM with parameters $\theta$. We perform a gradient descent step with training sample $d$ with learning rate $\eta$ such that $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Then,
@@
[…]
@@
[…]
@@
Lemma˜6 Define $\alpha_{j}^{\boldsymbol{x}}$ as